UT 20150330 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, WSeifert, APramskiy, DThompson (LBTO)
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator: Sallenson
LBTO Support: TSargent

Plan for the night

- do some AO-tests (longslit acquisition) weather permitting (forecast is not very promising)

- otherwise close-dome tests, this includes installationj of new LUCi SW version, active flexure compensation, repro-tests etc.

Summary

  • Successfull AO-present & offsets via scripts from LUCI

  • First AO spectrum

  • Power outage

Preparations

Tele/Skype-Con with LBTO-management, LBT daytime crew, Arectri- & LUCI-teams in the afternoon about current situation
Here is a summary of the discussion provided by AConrad:

=> LUC2/AGw will be brought down to the lab on Monday to fix the AGw-issue

=> If all works, we will go back on sky Wednesday night

=> The two nights w/o LUCi2/AGw will/can be used for engineering purposes

=> Possibly one or two of the engineering nights April 3-5 may used for continuation of AO-com (also for training purposes)

The Arcetri-people inspected the AGw/rotator after the meeting. The issues with the AO/AGw is not
the cable wrap but rather a board inside the AGw.

Details

All times are UT
Files are luci2.20150330.NNNN.fits

01:40 Steve waits a bit to open the dome (doesn't look tooooo bad actually), but one fat cloud is above us
Weather: sky partly clear, T = 6.5deg, low wind, humidity = 25%

03:42 Detector-tests for stepfunction:

clear/Ks, 2.73sec

LIR (idle-break): DIT=1, Nexp=20, #13-32
DIT=20, Nexp=1, #33
DIT=1, NDIT=20, #34-53
LIR(idle-wait): DIT=1, Nexp=20, #54-73
DIT=1, NDIT=20, #74-94
LIR (idle-break): DIT=1, NDIT=5, 6 times, #97-121

Reason for stepfunction is possibly persistence and the way idle-break works

04:50 Steve opens dome to do some testing

05:15 A bit more clearer now,

05:53 AO preset to A45_R8_0035 using script!
Failed a couple of times, finally worked, scr #125, slitview #126-128,
problems with Aladin at obs4 (different libraries than on LUCI-machine)
did alignment from Alex-laptop (worked), clouds, loop open...

07:23 Again AO preset to A45_R8_0035 using a script
scr #129, slitview #131, acq. worked, loop closed, offsets OK!

08:09 Again AO preset, trying this time with LS-acq., scr #139, slitview #140,
first AO-spectrum with G40 (#142) !!! Width of spectrum 0.15" fwhm (without correction for NCPA)!

08:26 A45_R8_0035, dice-5 pattern, repeated 10 times, check for accuracy of offsets

08:39 Restart, this time with absolute offsets in script, works but data need to be checked, #151-171

09:02 Restart, this time with field stop mask in FPU, #172-212

The Arcetir group wents to sleep, we are on our own with the AO

09:48 TYC 3058-268-1, double star with 6" separation,

09:52 Heavy power outage, its all dark

11:00 Back to life, running on generators now

11:20 Several attempts to start an AO-imaging script, the AO-system apparently didn't like the outage

11:45 Due to that LUCI failed also.

12:00 Overcast, end of night
Topic revision: r6 - 04 Apr 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback